OpenStudy (anonymous):

How can you use the subsitution method to solve the system x+3y=-3 Over 4x+3y=6

OpenStudy (anonymous):

I have to show the steps for the equation using the subsitution method for x+3y=-3 4x+3y=6

OpenStudy (texaschic101):

x + 3y = - 3 ---> x = -3y - 3 now sub - 3y - 3 in for x in the second equation to solve for y.

OpenStudy (anonymous):

Thanks but how would it be written?

OpenStudy (texaschic101):

One question .....is this the problem....x + 3y = - 3/4x + 3y = 6....you said " over " or is it : x + 3y = - 3 4x + 3y = 6

OpenStudy (anonymous):

It's what u wrote last I can't type it like this on my phone

OpenStudy (texaschic101):

ok..... x + 3y = -3 --> x = - 3y - 3 now sub -3y - 3 in for x 4x + 3y = 6 4(-3y - 3) + 3y = 6 (now distribute the 4 through the parenthesis and solve for y. Do you know how to do that ?

OpenStudy (anonymous):

Yep thank u so much :)

OpenStudy (texaschic101):

your welcome :)